OpenStudy (anonymous):

Find exact value of each trigonometric function cos 7π/3

OpenStudy (anonymous):

\[\cos \frac{ 7 π}{ 3 }\]

OpenStudy (freckles):

\[\text{ two hints } \\ \text{hint 1:} \frac{2\pi}{3}+\frac{5\pi}{3}=\frac{7\pi}{3} \\ \\ \text{ hint 2: } \cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b)\]

OpenStudy (anonymous):

\[\frac{ \sqrt{3}}{ 2 }*2\]

OpenStudy (freckles):

how did you get that?

OpenStudy (freckles):

\[\cos(\frac{7\pi}{3})= \cos(\frac{2\pi}{3}+\frac{5\pi}{3}) \\ =\cos(\frac{2\pi}{3})\cos(\frac{5\pi}{3})-\sin(\frac{2\pi}{3}) \sin(\frac{5\pi}{3}) \\ =?\] use unit circle to evaluate cos(2pi/3) and cos(5pi/3) and sin(2pi/3) and sin(5pi/3)

OpenStudy (anonymous):

because 2π=6π/3 then i added π/3

OpenStudy (freckles):

lol I should have thought about subtracting 2pi you're so brilliant

OpenStudy (freckles):

that does make things easier

OpenStudy (freckles):

\[\cos(\frac{7\pi}{3})=\cos(\frac{7\pi}{3}-2\pi)=\cos(\frac{\pi}{3})\]

OpenStudy (freckles):

and cos(pi/3) should actually be ...

OpenStudy (freckles):

look at the x-coordinate at pi/3

OpenStudy (anonymous):

would it be \[\sqrt{3}\]

OpenStudy (freckles):

cos(pi/3) find pi/3 on the unit circle and use the x-coordinate since we are dealing with cosine cos(pi/3)=1/2

OpenStudy (freckles):

so cos(7pi/3)=1/2

OpenStudy (anonymous):

\[\frac{ \sqrt{3} }{ 2 } *2=\sqrt{3}\]

OpenStudy (freckles):

I'm confused. Are you not wanting to evaluate cos(7pi/3) anymore?

OpenStudy (freckles):

cos function has period 2pi so cos(x)=cos(x-2pi)=cos(x+2pi)=cos(x+2pi*n) where n can be any integer cos(7pi/3)=cos(7pi/3-2pi)=cos(pi/3)=1/2 we just look at the x-coordinate at 7pi/3

OpenStudy (anonymous):

i think i got it now thanx for your help

OpenStudy (freckles):

*we just look at the x-coordinate at pi/3

OpenStudy (freckles):

another example: \[\cos(\frac{19\pi}{6})=\cos(\frac{19\pi}{6}-2\pi)=\cos(\frac{7\pi}{6})=\frac{-\sqrt{3}}{2}\]